# Belief Revision: The Adaptability of Large Language Models Reasoning

### v1

@@ Abstract
The capability to reason from text is crucial for real-world NLP applications. Real-world scenarios often involve incomplete or evolving data. In response, individuals update their beliefs and understandings accordingly. However, most existing evaluations assume that language models (LMs) operate with consistent information. We introduce Belief-R111 We will release the dataset and code upon acceptance., a new dataset designed to test LMs’ belief revision ability when presented with new evidence. Inspired by how humans suppress prior inferences, this task assesses LMs within the newly proposed delta reasoning ($\Delta R$) framework. Belief-R features sequences of premises designed to simulate scenarios where additional information could necessitate prior conclusions drawn by LMs. We evaluate $\sim$30 LMs across diverse prompting strategies and found that LMs generally struggle to appropriately revise their beliefs in response to new information. Further, models adept at updating often underperformed in scenarios without necessary updates, highlighting a critical trade-off. These insights underscore the importance of improving LMs’ adaptiveness to changing information, a step toward more reliable AI systems.
@@ 1 Introduction
Through Belief-R, we evaluate the belief revision ability of small and large scale LMs using different prompting techniques. Our study shows that these models often fail to adjust their responses when presented with new information that necessitates adjustments. We further reveal a critical limitation: they confront a performance trade-off between updating and maintaining their prior beliefs. Models that perform better in the cases where an update is needed, typically faltered on the other. Furthermore, better prompting methods also fail to significantly enhance this capability. These insights underscore a need for strategies to enhance model’s capability to correctly update or maintain its initial beliefs when faced with new evidence to ensure its reliability across evolving scenarios.
@@ 2 Related Works
Reasoning is one of the fundamental intelligent behaviors, essential for solving complex real-world tasks Huang and Chang (2023). Many studies test this behaviour in LMs by setting up various tasks. For instance, some create simple tasks to check if a system can answer questions by connecting facts or using basic logic Weston et al. (2016). Others design more advanced tests to evaluate inductive, deductive, and abductive reasoning Sinha et al. (2019); Saparov et al. (2024); Bhagavatula et al. (2020). Some benchmarks replicate real-world complexities by presenting partial or conflicting informations Talmor et al. (2020); Arabshahi et al. (2021); Sprague et al. (2022); Han et al. (2024); Kazemi et al. (2024). Our research pushes this boundary further, focusing on scenarios where information evolves, presenting queries that neccessitate a dynamic update of prior beliefs in light of this new evidence. We note the comparison in Table 1.
@@ 3.1 Background and notation
For set of query sentences $\chi$, it encompasses a set of premises $\Gamma{=}\{\gamma_{1},\dots,\gamma_{N}\}$ that could imply a set of conclusions $\Phi{=}\{\varphi_{1},\dots,\varphi_{M}\}$. We denote reasoner’s belief set as a set of sentences $\mathcal{B}$ to represent a contextually fixed background knowledge of $\chi$. In this regard, $\mathcal{B}$ is a tuple that contains set of premises and conclusions: $\mathcal{B}{=}(\Gamma,\Phi)$. In presence of new information $\gamma_{N\texttt{+}1}$, the belief revision concept allow us to infer conclusion $\varphi_{M\texttt{+}1}$ if it is rational to believe $\varphi_{M\texttt{+}1}$ after acknowledging $\gamma_{N\texttt{+}1}$.
@@ 4.1.2 Ground-truth formulation
To further validate the implied commonsense interaction of the third premises, whether it serves as alternative or additional condition, we manually annotate the final conclusions through a crowdsource annotation task at Appenfootnote 2 (see Appendix B). We cater the variability arises from different interpretations from diverse human readers by asking 5 workers to annotate each problem and then take the majority voting out of them to set the agreed options as the ground truths. Upon further inspection, we found some annotations that logically invalid, i.e. answering $\neg q$ in questions with modus ponens inferences or answering $p$ in modus tollens inferences. We view such cases as non modus ponens (or tollens) inferences and specifically treat the annotation similarly with answering c) $\Diamond q\land\Diamond\neg q$.
@@ 4.2 Quality check
Throughout the data construction phase, we assign one expert to review of the logical formations to ensure they follow the intended structure. We also further gauge the quality of the generated data by reviewing 100 randomly chosen samples to confirm on the context and logical consistency. We conducted a human evaluation via Appen [^2] https://appen.com/, with three native English speakers assessing each sample’s quality. They unanimously confirmed that the conditional relationships in the premises were logically sound across all samples, i.e.  that $q$ entails $p$ and $q$ entails $r$ in both of the conditional premises. We also attach the annotation guidelines in Appendix B.
@@
To enhance the quality of our dataset for more reliable evaluation, we refined it by focusing on consensus among annotators. For each question, we utilize the answers manually labeled by five independent workers. We retained only the questions with strong majority agreement (at least four out of five annotators concurred). This filtering retained $\sim$65of the original data.
@@ 4.3 Statistics of Belief-R
Figure 4: BREU score evaluation on belief revision capabilities in Belief-R on various models sorted by the BREU score. While larger-scale LLMs tend to achieve higher BREU score, the performance is far lower compared their basic logical inference at $t$, showcasing limited capability of LLMs in performing belief revision.
@@ 5 Experiment Settings
The primary goal of our experiments is to investigate whether LMs possess the capability to perform belief revision in their reasoning processes. We report the accuracies in the Belief Update (BU-Acc) and the Belief Maintain (BM-Acc) subsets to indicate LM capabilities in updating and maintaining its beliefs in which it has to do so. We further introduce a novel metric, BREU (Belief Revision Evaluation Understudy), to assess LMs’ belief revision ability, by averaging BU-Acc and BM-Acc equally. The goal of BREU is to gauge whether the model accurately decides when to update or maintain its prior beliefs. We then benchmark publicly-available LMs and design series of experiments through $\mathbf{\Delta R}$ framework. We perform zero-shot-classification on series of smaller to larger scales pre-trained and finetuned LMs, and prompt LLMs generations through API.
@@ 6 Result and Analysis
(a) Modus Ponens / Modus Tollens
(b) Events / Mental States
(c) Prompting Methods
@@ 7 Discussion
Belief update on mental states effect entities is more challenging than events.
@@
Better prompting methods yield limited gain on belief revision.
We explore how different prompting methods affect belief revision abilities of LLMs on average. Figure 5(c) shows that CoT, which encourages LLMs to elicit reasoning steps, does not significantly enhance belief revision. While this may stem from its vulnerability to missing-step errors Wang et al. (2023b), attempts to correct these errors with the PS prompting offer minimal benefits, improving only by $\sim$1of BREU. This suggests the ability to revise beliefs could still be absent despite elicitation of reasoning steps.
@@ 8 Conclusion
The ability to reason and adapt to changing information is crucial for NLP applications in the real world. Most evaluations assume static knowledge environment, which does not prepare models for dynamic real-life scenarios. To address this, Belief-R is introduced as a diagnostic dataset for evaluating belief revision capability in LMs. Through Belief-R and a novel evaluation framework for evaluating reasoning in a dynamically evolving environment, $\mathbf{\Delta R}$, we reveal that current models struggle with updating their beliefs in response to new information, highlighting the need for improved adaptability and reliability. Our work emphasizes the significance of enhancing the capability of AI models to reason with evolving data for real-world readiness.
@@
Our study on belief revision using the Belief-R dataset via the $\mathbf{\Delta R}$ framework focuses on belief changes driven by logical inferences like modus ponens and modus tollens, which may not fully represent the complexity of real-world belief revision that often includes a broader range of scenarios and subtleties. Moreover, our methodology primarily considers the introduction of new premises as the trigger for belief revision, overlooking how beliefs might change through re-evaluation of existing knowledge or shifts in perspective in the absence of new information (i.e. in Kronemyer and Bystritsky (2014)). Additionally, our approach to simulate future data is constrained by our inability to determine what LMs have previously known and by resource limitations that restrict the training of large-scale models from scratch.
@@
This research explores how well LMs can revise their beliefs when faced with new information, which is crucial for their use in constantly changing real-world situations. We created a reasoning evaluation dataset to test whether LMs can revise their beliefs correctly or if they stick to their initial assumptions. This is important for using LMs in areas where being accurate and up-to-date is vital, like healthcare or legal advice. In example, being able to revise beliefs appropriately could help prevent LMs from repeating outdated or wrong information, making them more reliable and trustworthy. Plus, LMs that can refresh their understanding according to new societal norms can avoid perpetuating biases, contributing to the fair and ethical use of AI. We consider this a promising and significant area for research. We construct the dataset using events, causes, and effects from ATOMIC and the construction template is designed and reviewed manually and attached in this paper. We utilized crowd-sourced annotators who voluntarily participated through the platform Appenfootnote 2, choosing tasks they deemed fairly compensated. The annotators were presented with multiple-choice tasks predefined to avoid bias and protect privacy, ensuring an ethical annotation process.
@@ C.1 LLMs logical reasoning ability are not robust in the presence of distractors
C.2 Better prompting methods yield limited gain on belief revision.

### v2

@@ Abstract
The capability to reason from text is crucial for real-world NLP applications. Real-world scenarios often involve incomplete or evolving data. In response, individuals update their beliefs and understandings accordingly. However, most existing evaluations assume that language models (LMs) operate with consistent information. We introduce Belief-R111 The code and dataset are available at https://github.com/HLTCHKUST/belief-revision, a new dataset designed to test LMs’ belief revision ability when presented with new evidence. Inspired by how humans suppress prior inferences, this task assesses LMs within the newly proposed delta reasoning ($\Delta R$) framework. Belief-R features sequences of premises designed to simulate scenarios where additional information could necessitate prior conclusions drawn by LMs. We evaluate $\sim$30 LMs across diverse prompting strategies and found that LMs generally struggle to appropriately revise their beliefs in response to new information. Further, models adept at updating often underperformed in scenarios without necessary updates, highlighting a critical trade-off. These insights underscore the importance of improving LMs’ adaptiveness to changing information, a step toward more reliable AI systems.
@@ 1 Introduction
Through Belief-R, we evaluate the belief revision ability of small and large scale LMs using different prompting techniques. Our study shows that these models are incapable of revising their prior beliefs. We further reveal a critical limitation: they confront a performance trade-off between updating and maintaining their prior beliefs. Models that perform better in the cases where an update is needed, typically faltered on the other. Furthermore, better prompting methods also fail to significantly enhance this capability. These insights underscore a need for strategies to enhance model’s capability to correctly update or maintain its initial beliefs when faced with new evidence to ensure its reliability across evolving scenarios.
@@ 2 Related Works
Reasoning is one of the fundamental intelligent behaviors, essential for solving complex real-world tasks Huang and Chang (2023). One works test this behaviour by creating simple tasks to comprehensively check if a system can answer questions by connecting facts or using basic logic Weston et al. (2016). Others design more advanced tests to evaluate inductive, deductive, and abductive reasoning Sinha et al. (2019); Saparov et al. (2024); Bhagavatula et al. (2020). Some benchmarks replicate real-world complexities by presenting partial or conflicting informations Arabshahi et al. (2021); Sprague et al. (2022); Han et al. (2024); Kazemi et al. (2024). Belief revision focuses on the adaptability problem: whether the model properly revises prior beliefs as new information emerges. We extend the reasoning evaluation by focusing on scenarios where information evolves, presenting queries that require dynamic updates of prior beliefs in light of new evidence. This is distinct from other existing reasoning tasks involving incomplete or contradictory information, i.e. in Talmor et al. (2020); Bhagavatula et al. (2020); Kazemi et al. (2024), since they assume a static environment and focus on filling the gap or resolving the contradiction. We further note the comparison in Table 1.
@@ 3.1 Background and notation
For set of query sentences $\chi$, it encompasses a set of premises $\Gamma{=}\{\gamma_{1},\dots,\gamma_{N}\}$ that could imply a set of conclusions $\Phi{=}\{\varphi_{1},\dots,\varphi_{M}\}$. In this work, we conceptualize "belief" similarly to its usage in dialogue systems, where it represents what the system currently considers true based on the context Feng et al. (2023); van Niekerk et al. (2020). We denote reasoner’s belief set as a set of sentences $\mathcal{B}$ to represent a contextually fixed background knowledge of $\chi$. In this regard, $\mathcal{B}$ is a tuple that contains set of premises and conclusions: $\mathcal{B}{=}(\Gamma,\Phi)$. In presence of new information $\gamma_{N\texttt{+}1}$, the belief revision concept allow us to infer conclusion $\varphi_{M\texttt{+}1}$ if it is rational to believe $\varphi_{M\texttt{+}1}$ after acknowledging $\gamma_{N\texttt{+}1}$.
@@ 4.1.2 Ground-truth formulation
To further validate the implied commonsense interaction of the third premises, whether it serves as alternative or additional condition, we manually annotate the final conclusions through a crowdsource annotation task at Appensection 4.3 (see Appendix B). We cater the variability arises from different interpretations from diverse human readers by asking 5 workers to annotate each problem and then take the majority voting out of them to set the agreed options as the ground truths. Upon further inspection, we found some annotations that logically invalid, i.e. answering $\neg q$ in questions with modus ponens inferences or answering $p$ in modus tollens inferences. We view such cases as non modus ponens (or tollens) inferences and specifically treat the annotation similarly with answering c) $\Diamond q\land\Diamond\neg q$.
@@ 4.2 Quality check
Throughout the data construction phase, we assign one expert to review of the logical formations to ensure they follow the intended structure. We also further gauge the quality of the generated data by reviewing 100 randomly chosen samples to confirm on the context and logical consistency. We conducted a human evaluation via Appen22footnotemark: 2, with three native English speakers assessing each sample’s quality. They unanimously confirmed that the conditional relationships in the premises were logically sound across all samples, i.e.  that $q$ entails $p$ and $q$ entails $r$ in both of the conditional premises. We also attach the annotation guidelines in Appendix B.
@@
To enhance the quality of our dataset for more reliable evaluation, we refine it by focusing on consensus among annotators. For each question, we utilize answers manually labeled by five independent workers. We measured inter-annotator agreement using Gwet’s AC1 Gwet (2008) [^3] https://pypi.org/project/irrCAC/ since it is better at handling high agreement scenarios, with criteria used in Wongpakaran et al. (2013) to interpret the level of agreement.
We initially measured the annotation agreement which yielded a moderate score of 0.573. Further, we observe that some beliefs are naturally more subjective than others. To make the evaluation dataset more rigorous, we only take the ones with the higher agreement, as commonly practiced in subjective tasks such as sentiment analysis Bobicev and Sokolova (2017). We retain only questions with strong majority agreement (at least 4 out of 5 annotators concurred). Post-filtering, as we retain $\sim$65of the original data, the score improved to 0.697 and indicated a substantial agreement.
Figure 4: BREU score evaluation on belief revision capabilities in Belief-R on various models sorted by the BREU score. While larger-scale LLMs tend to achieve higher BREU score, the performance is far below their basic logical inferences at $t$ (Acc@t), showcasing limited capability of LLMs in performing belief revision.
@@ 4.3 Statistics of Belief-R
33footnotetext: https://appen.com/
@@ 5 Experiment Settings
The primary goal of our experiments is to investigate whether LMs possess the capability to perform belief revision in their reasoning processes. Concretely, we consider the models’ predictions on the Basic @t category in the evaluation dataset as the models’ initial belief. Thus, changes in accuracy on the Belief Update and Belief Maintain categories directly reflects belief revision. We report accuracies in the Belief Update (BU-Acc) and the Belief Maintain (BM-Acc) subsets to indicate LMs’ capabilities in updating and maintaining their beliefs when required. We further introduce a novel metric, BREU (Belief Revision Evaluation Understudy), to assess LMs’ belief revision ability, by averaging BU-Acc and BM-Acc equally. The goal of BREU is to gauge whether the model accurately decides when to update or maintain its prior beliefs. We then benchmark publicly-available LMs and design series of experiments through $\mathbf{\Delta R}$ framework. We perform zero-shot-classification on series of smaller to larger scales pre-trained and finetuned LMs, and prompt LLMs generations through API.
(a) Modus Ponens / Modus Tollens
(b) Events / Mental States
(c) Prompting Methods
@@ 7 Discussion
Belief update on abstract concept is more challenging for LLMs.
@@
Better prompting method does not help on belief revision.
We explore how different prompting methods affect belief revision abilities of LLMs on average. Figure 5(c) shows that CoT, which encourages LLMs to elicit reasoning steps, does not significantly enhance belief revision. While this may stem from its vulnerability to missing-step errors Wang et al. (2023b), attempts to correct these errors with the PS prompting offer minimal benefits, improving only by $\sim$1of BREU. This suggests the ability to revise beliefs could still be absent despite elicitation of reasoning steps. We put a more detailed analysis in Appendix C.2.
@@ 8 Conclusion
The ability to reason and adapt to changing information is crucial for NLP applications in the real world. Most evaluations assume static knowledge environment, which does not prepare models for dynamic real-life scenarios. To address this, Belief-R is introduced as a diagnostic dataset for evaluating belief revision capability in LMs. Through Belief-R and a novel evaluation framework for evaluating reasoning in a dynamically evolving environment, $\mathbf{\Delta R}$, we reveal that current models struggle with updating their beliefs in response to new information, highlighting the need for improved adaptability and reliability. By exposing these limitations, our work underscores the imperative of developing AI models capable of reasoning adeptly with evolving data, thereby propelling them closer to real-world applicability and robustness.
9 Future Work
Observing the limited gain offered by the current prompting strategy, we suggest that future developments aim to enhance LMs’ understanding and handling of premise dependencies. A particularly promising direction is to integrate LMs with deterministic symbolic solvers to convert problems into logical rules through chain-of-thought reasoning to aim for a deeper comprehension of the intricate dependencies among different premises. Future work could aim to overcome current limitations and pave the way for AI systems to be reliable across evolving scenarios.
Acknowledgements
We thank Yejin Bang for the insightful discussions. This work has been partially funded by the PF20-43679 Hong Kong PhD Fellowship Scheme and the Hong Kong Fellowship Scheme by the Hong Kong Research Grants Council (RGC), and Care-E Project (FS116).
@@ Acknowledgements
Our study on belief revision using the Belief-R dataset via the $\mathbf{\Delta R}$ framework focuses on belief changes driven by logical inferences like modus ponens and modus tollens, which may not fully represent the complexity of real-world belief revision that often includes a broader range of scenarios and subtleties. To add, our methodology primarily considers the introduction of new premises as the trigger for belief revision, overlooking how beliefs might change through re-evaluation of existing knowledge or shifts in perspective in the absence of new information (i.e. in Kronemyer and Bystritsky (2014)).
Furthermore, how to define models’ beliefs is a debatable question and there are some works inspecting models’ internal beliefs by probing the hidden states Burns et al. ; Zou et al. (2023). n our work, however, we conceptualize "belief" similarly to its usage in dialogue systems, where it represents what the system currently considers true based on the context Feng et al. (2023); van Niekerk et al. (2020). Further exploration on examining the belief through other approaches (e.g., probing hidden states of LLMs) lies outside this paper’s scope and we leave it as future work.
@@
This research explores how well LMs can revise their beliefs when faced with new information, which is crucial for their use in constantly changing real-world situations. We created a reasoning evaluation dataset to test whether LMs can revise their beliefs correctly or if they stick to their initial assumptions. This is important for using LMs in areas where being accurate and up-to-date is vital, like healthcare or legal advice. In example, being able to revise beliefs appropriately could help prevent LMs from repeating outdated or wrong information, making them more reliable and trustworthy. Plus, LMs that can refresh their understanding according to new societal norms can avoid perpetuating biases, contributing to the fair and ethical use of AI. We consider this a promising and significant area for research. We construct the dataset using events, causes, and effects from ATOMIC and the construction template is designed and reviewed manually and attached in this paper. We utilized crowd-sourced annotators who voluntarily participated through the platform Appensection 4.3, choosing tasks they deemed fairly compensated. The annotators were presented with multiple-choice tasks predefined to avoid bias and protect privacy, ensuring an ethical annotation process.
@@ C.1 LLMs logical reasoning ability are not robust in the presence of distractors
C.2 Details on prompting methods variations on their gain at belief revision.